If 2x + 5 + -25 and -3m -6 = 48, what is the product of x and m?
Answer by Boreal(15235)

2x + 5 =-25, I put the equal sign in. 2x=-30, x=-15
-3m -6 = 48, -3m=54, m=-18
Their product is +270.